# Highly Efficient Rank-Adaptive Sweep-based SI-DSA for the Radiative Transfer Equation via Mild Space Augmentation

### 26 Mar 2026 v1

@@ 3.4 Incremental basis and reduced operator update
In the inner-loop iteration of our low-rank SI, the space of spatial modes is progressively enlarged by adding snapshots of angular fluxes for sampled angles. By exploiting this hierarchical structure, the spatial basis ${\bf{X}}^{(n,k)}$ and Galerkin projected operators onto its span can be updated incrementally. Specifically, the orthonormal spatial basis ${\bf{X}}^{(n,k)}$ can be updated through MGS-RO in Alg. 2. Reorthogonalization is selectively applied to mitigate the loss of orthogonality in ${\bf{X}}^{(n,k)}$ caused by numerical error accumulation in the MGS process.
Algorithm 2 Modified Gram–Schmidt with selective reorthogonalization (MGS-RO)
@@
4:  Compute $\hat{\mathbf{c}}={\bf{X}}^{T}\bm{\psi}^{(n)}_{s^{(k)}_{l}}$ and $\hat{\bm{\psi}}=\bm{\psi}^{(n)}_{s^{(k)}_{l}}-{\bf{X}}\hat{\mathbf{c}}$.
5:  Define $\tilde{\mathbf{c}}=\bigl(\hat{\mathbf{c}}^{T},\ \|\hat{\bm{\psi}}\|_{2}\bigr)^{T}$.
6:  Update ${\bf{X}}:=\begin{bmatrix}{\bf{X}}&\hat{\bm{\psi}}/\|\hat{\bm{\psi}}\|_{2}\end{bmatrix}$ and $\mathbf{R}:=\begin{bmatrix}\mathbf{R}&\tilde{\mathbf{c}}\end{bmatrix}$.
7:end for
@@
9:  Recompute a QR factorization:
${\bf{X}}^{(n,k)}\mathbf{R}^{(n,k)}=\mathrm{qr}\!\left(\begin{bmatrix}{\bf{X}}^{(n,k-1)}&\bm{\psi}^{(n)}_{s^{(k)}_{1}}&\dots&\bm{\psi}^{(n)}_{s^{(k)}_{p}}\end{bmatrix}\right)$.
@@ 4 Numerical examples
Before delving into details, we outline numerical tests performed and their purposes. In Sec. 4.1, for a homogeneous medium ranging from weak to strong scattering, we perform a refinement study to investigate the influence of mesh resolution and the effectiveness of our method in different regimes. In Sec. 4.2, we simulate a multiscale variable scattering problem and investigate the influence of $p$ and $q$ in our sampling step. In Sec. 4.3, a pin-cell problem with strong material discontinuities is solved repeatedly to demonstrate the robustness of the method with respect to randomness in the subsampling step. In Sec. 4.4, a lattice problem is solved using both the proposed method and a modified version to illustrate the necessity of the incremental update of the spatial basis and projection operators.
@@ 4.1 Homogeneous medium in different regime
22 $56$ $3.62\times 10^{-8}$ $4.28\times 10^{-9}$ $45.13\%$ 55 55 $1.70\times$
33 $93$ $1.38\times 10^{-7}$ $1.07\times 10^{-8}$ $33.30\%$ 55 55 $2.77\times$
44 $127$ $2.31\times 10^{-7}$ $7.99\times 10^{-9}$ $25.58\%$ 55 55 $3.24\times$
55 $160$ $6.60\times 10^{-7}$ $1.41\times 10^{-8}$ $20.63\%$ 55 55 $4.86\times$
@@
22 $64$ $2.31\times 10^{-9}$ $1.22\times 10^{-9}$ $51.57\%$ 88 88 $1.37\times$
33 $123$ $2.84\times 10^{-7}$ $1.84\times 10^{-8}$ $44.05\%$ 88 88 $1.86\times$
44 $175$ $1.08\times 10^{-7}$ $7.31\times 10^{-9}$ $35.25\%$ 88 88 $2.29\times$
55 $234$ $6.70\times 10^{-7}$ $1.57\times 10^{-8}$ $30.17\%$ 88 88 $2.44\times$
@@
22 $64$ $3.23\times 10^{-7}$ $2.96\times 10^{-7}$ $51.57\%$ 88 88 $1.55\times$
33 $141$ $1.28\times 10^{-7}$ $5.49\times 10^{-8}$ $50.49\%$ 88 88 $1.04\times$
44 $194$ $6.59\times 10^{-8}$ $6.07\times 10^{-9}$ $39.08\%$ 88 88 $1.57\times$
55 $229$ $7.05\times 10^{-8}$ $4.37\times 10^{-9}$ $29.52\%$ 88 88 $1.64\times$
@@
22 $53$ $7.11\times 10^{-8}$ $7.06\times 10^{-8}$ $2.71\% 44$ 44 $1.76\times$
33 $61$ $1.80\times 10^{-7}$ $1.79\times 10^{-7}$ $21.84\%$ 44 44 $3.73\times$
44 $63$ $1.71\times 10^{-7}$ $1.70\times 10^{-7}$ $12.69\%$ 44 44 $6.71\times$
55 $63$ $1.93\times 10^{-7}$ $1.91\times 10^{-7}$ $8.12\%$ 55 55 $13.88\times$
@@ 4.2 Multiscale variable scattering problem
Robust accuracy and convergence against various $p$ and $q$. Across all choices of $p$ and $q$, our low-rank SI-DSA converges in the same number of iterations as the full-rank SI-DSA. Moreover, the difference between the scalar fluxes produced by the low-rank and full-rank methods, $\|\bm{\phi}_{\textrm{LR}}-\bm{\phi}_{\textrm{FR}}\|_{2}$ is on the order of $10^{-7}$-$10^{-8}$, at least one order of magnitude smaller than the threshold $\epsilon_{\textrm{SI-SA}}=10^{-6}$ used in the stopping criterion of SI-DSA, i.e. $\|\bm{\phi}^{(n)}-\bm{\phi}^{(n-1)}\|_{\infty}\leq\epsilon_{\textrm{SI-SA}}$ .
Robustness against $p$ and $q$. We observe that, for various $p$ and $q$, the low-rank SI-DSA consistently reduces the DOFs required to store the angular flux by approximately $38\%$. Moreover, the reduction in computational time is around $1.20\times$ for most cases, except for $1.05\times$ with $(p,q)=(1,16)$. The slightly smaller speedup in this case may be attributed to the increased computational cost in the sampling step.
Mild space augmentation. In Fig. 3, we present the effective rank and the oversampling ratio during low-rank SI for various $p$ and $q$. With $p=1$, the oversampling ratio remains below $5\%$ for all choices of $q$. For fixed $q=8$, increasing the number of sampled angles per iteration leads to a higher oversampling ratio, but it never exceeds $20\%$. These observations verifies the effectiveness of the proposed incremental strategy in preventing overly aggressive space augmentation during rank adaptation. We emphasize that a rank adaptive method using the standard $O(r)$ space augmentation may even become slower than the full-rank solver, since the solution rank is close to $38\%$ of the full rank under the present setup.
@@
44 $281$ $3.90\times 10^{-6}$ $1.06\times 10^{-7}$ $36.22\%$ 77 77 $1.17\times$
88 $298$ $6.29\times 10^{-7}$ $8.87\times 10^{-8}$ $38.42\%$ 77 77 $1.19\times$
$12$ $297$ $29\times 10^{-7} 1.1$$14\times 10^{-7}$ $38.29\%$ 77 77 $1.22\times$
$16$ $299$ $5.49\times 10^{-8}$ $3.58\times 10^{-8}$ $38.54\%$ 77 77 $1.05\times$
@@
11 $298$ $6.29\times 10^{-7}$ $8.87\times 10^{-8}$ $38.42\%$ 77 77 $1.19\times$
22 $298$ $3.53\times 10^{-6}$ $1.08\times 10^{-7}$ $38.42\%$ 77 77 $1.29\times$
33 $315$ $2.02\times 10^{-6}$ $2.89\times 10^{-7}$ $40.61\%$ 77 77 $1.18\times$
@@ 4.3 Pin-cell problem
Effectiveness and robustness with respect to the randomness in the algorithm. From the accuracy perspective, the difference between the solutions produced by the low-rank SI-DSA and the full-rank SI-DSA is consistently much smaller than $\epsilon_{\textrm{SI-SA}}=10^{-6}$. In $7$ out of $8$ runs, our low-rank SI-DSA converges with exactly the same number of iterations as its full-rank counterpart, while one run requires one more iteration.
The solution ranks for all runs of the low-rank SI-DSA are close to $42\%$ of the full rank. Moreover, even in this challenging scenario, thanks to the efficiency of the proposed sampling strategy and the mild space augmentation in the rank-adaptation step, the method achieves at least a $1.11\times$ speedup and an average speedup of $1.17\times$. In the best run, a speedup of approximately $1.40\times$ is obtained.
Figure 5 shows the convergence history and the oversampling ratio during the low-rank SI iterations for the last run. Except for the second-to-last iteration, the change of the scalar flux in the $k$-th iteration, $\|\bm{\phi}^{(k)}-\bm{\phi}^{(k-1)}\|_{2}$, for the low-rank and full-rank SI-DSA almost coincide. The oversampling ratio remains below 9% throughout all iterations. The observations for the other runs are similar and are therefore omitted.
Finally, we compute the effective rank of the full-rank angular flux by applying truncated SVD to its matrix representation with the same truncation tolerance as in the low-rank method. The effective rank of the full-rank solution is 222, while the ranks of the low-rank solutions range from 207 to 218, with an average of 212.63.
@@
Figure 5: Convergence history, effective rank and oversampling ratio during the low-rank SI for the pin-cell problem in Sec. 4.3. Left: convergence history. Right: oversampling ratio.
@@
Solution rank $212.63$ ($\pm 3.00$) $207$ $218$
$\|\psi_{\text{LR}}-\psi_{\text{FR}}\|_{2}$ $2.08(\pm 1.73)\times 10^{-7}$ $9.47\times 10^{-8}$ $6.53\times 10^{-7}$
$\|\phi_{\text{LR}}-\phi_{\text{FR}}\|_{2}$ $1.14(\pm 1.49)\times 10^{-7}$ $2.29\times 10^{-8}$ $5.05\times 10^{-7}$
@@
4.4 Lattice problem
Our final test is a lattice problem with zero inflow boundary conditions, and its configuration is shown in Fig. 6. The black region corresponds to pure absorption with $(\sigma_{a},\sigma_{s})=(100,0)$, while the white and orange regions correspond to pure scattering with $(\sigma_{a},\sigma_{s})=(0,1)$. A constant source $G(x,y)=1$ is imposed in the orange region. The computational domain $0,5]^{2}$ is discretized using an $80\times 80$ mesh in physical space and the CL(32,16) quadrature in angular space. The scalar fluxes obtained by the low-rank and full-rank SI-DSA are shown on a logarithmic scale in Fig. 6 and match each other well.
This example is used to demonstrate the necessity of the incremental update of the spatial basis and the projected reduced operators. In particular, we compare the performance of our method with incremental updates against a modified version that replaces MGS-RO with direct QR decomposition in the basis update. We note that the incremental update of projection operators relies on the hierarchical structure of the spatial basis; therefore, incremental updates of the projected operators cannot be applied after this replacement.
Necessity of incremental update. As shown in the left plot of Fig. 7, the histories of $\|\bm{\phi}^{(k)}-\bm{\phi}^{(k-1)}\|_{2}$ for the low-rank and full-rank SI-DSA almost coincide. As reported in Tab. 4, the proposed low-rank SI-DSA achieves approximately $45.22\%$ compression in DOFs required to store the angular flux, converges in the same number of iterations as the full-rank SI-DSA, and achieves a $1.15\times$ speedup over it.
Although replacing the incremental update with direct QR decomposition does not affect the iteration count, the leading digits of the error compared with the full-rank solution and the memory savings remain essentially unchanged. However, the resulting low-rank solver becomes significantly slower, in fact approximately $1.53\times$ slower than the full-rank method.
Figure 7: Convergence history, effective rank and oversampling ratio during low-rank source iterations for the lattice problem in Sec. $4.$ Left: convergence history. Right: oversampling ratio.
MGS $227$ $4.63\times 10^{-7}$ $3.27\times 10^{-8}$ $45.22\%$ $11$ $11$ $1.15\times$
QR $227$ $4.63\times 10^{-7}$ $3.27\times 10^{-8}$ $45.22\%$ $11$ $11$ $0.65\times$
Table 4: Results for the lattice problem in Sec. $4.$ MGS: modified Gram-Schimdt with selective reorthogonalziation to update spatial basis and incremental update of projected operators. QR: using direct QR decomposition to update spatial basis. FR/LR-Iter: number of iterations required for convergence by the full/low-rank SI-DSA.

### 13 Apr 2026 v2

@@ 3.4 Incremental basis and reduced operator update
In the inner-loop iteration of our low-rank SI, the space of spatial modes is progressively enlarged by adding snapshots of angular fluxes for sampled angles. By exploiting this hierarchical structure, the spatial basis ${\bf{X}}^{(n,k)}$ and Galerkin projected operators onto its span can be updated incrementally. Specifically, the orthonormal spatial basis ${\bf{X}}^{(n,k)}$ can be updated via a variant of Modified Gram-Schmidt with double projection/one reorthognal pass (MGS2) [daniel1976reorthogonalization] in Alg. 2. It It is well-known that the orthonormal bases obtained by MGS may lose orthogonality quickly due to error accumulation [daniel1976reorthogonalization], and the reprojection step in Alg. 2 is introduced to mitigate this loss of orthogonality. Full QR pass is further selectively applied to re-orthogonalize basis ${\bf{X}}^{(n,k)}$ when it loses the orthogonality in ${\bf{X}}^{(n,k)}$. We highlight that, compared to MGS without reprojection, the MGS2 in Alg. 2 can substantially reduce the number of full QR reorthogoanlization in practice (See Sec. 4.2 for details).
Algorithm 2 Modified Gram–Schmidt with double projection and selective reorthogonalization (MGS2-RO)
@@
4:  First projection: $\hat{\mathbf{c}}={\bf{X}}^{T}\bm{\psi}^{(n)}_{s^{(k)}_{l}}$ and $\hat{\bm{\psi}}=\bm{\psi}^{(n)}_{s^{(k)}_{l}}-{\bf{X}}\hat{\mathbf{c}}$.
[…]
6:  Define $\tilde{\mathbf{c}}=\bigl(\hat{\mathbf{c}}^{T},\ \|\hat{\bm{\psi}}\|_{2}\bigr)^{T}$.
7:  Update ${\bf{X}}:=\begin{bmatrix}{\bf{X}}&\hat{\bm{\psi}}/\|\hat{\bm{\psi}}\|_{2}\end{bmatrix}$ and $\mathbf{R}:=\begin{bmatrix}\mathbf{R}&\tilde{\mathbf{c}}\end{bmatrix}$.
@@
9:   Reorthonalization through full QR:
${\bf{X}}\mathbf{R}^{(n,k)}=\mathrm{qr}\!\left(\begin{bmatrix}{\bf{X}}^{(n,k-1)}&\bm{\psi}^{(n)}_{s^{(k)}_{1}}&\dots&\bm{\psi}^{(n)}_{s^{(k)}_{p}}\end{bmatrix}\right)$.
@@ 4 Numerical examples
Before delving into details, we outline numerical tests performed and their purposes. In Sec. 4.1, for a homogeneous medium ranging from weak to strong scattering, we perform a refinement study to investigate the influence of mesh resolution and the effectiveness of our method in different regimes. In Sec. 4.2, a lattice problem is solved to illustrate the necessity of using MGS2 to incrementally update spatial basis. In Sec. 4.3, we simulate a multiscale variable scattering problem and investigate the influence of $p$ and $q$ in our sampling step. In Sec. 4.4, a pin-cell problem with strong material discontinuities is solved repeatedly to demonstrate the robustness of the method with respect to randomness in the subsampling step.
@@ 4.1 Homogeneous medium in different regime
22 $55$ $1.96\times 10^{-6}$ $1.55\times 10^{-7}$ $44.32\%$ 55 55 $1.50\times$
33 $93$ $1.38\times 10^{-7}$ $1.07\times 10^{-8}$ $33.30\%$ 55 55 $2.37\times$
44 $128$ $1.06\times 10^{-7}$ $3.65\times 10^{-9}$ $25.78\%$ 55 55 $2.77\times$
55 $160$ $6.60\times 10^{-7}$ $1.41\times 10^{-8}$ $20.63\%$ 55 55 $4.58\times$
@@
22 $64$ $2.41\times 10^{-10}$ $1.41\times 10^{-10}$ $51.57\%$ 88 88 $1.75\times$
33 $127$ $2.04\times 10^{-8}$ $2.47\times 10^{-9}$ $45.48\%$ 88 88 $1.90\times$
44 $175$ $1.08\times 10^{-7}$ $7.31\times 10^{-9}$ $35.25\%$ 88 88 $2.22\times$
55 $234$ $6.70\times 10^{-7}$ $1.57\times 10^{-8}$ $30.17\%$ 88 88 $2.67\times$
@@
22 $64$ $3.20\times 10^{-7}$ $2.94\times 10^{-7}$ $51.57\%$ 88 88 $1.59\times$
33 $143$ $8.12\times 10^{-8}$ $6.49\times 10^{-9}$ $51.21\%$ 88 88 $1.43\times$
44 $194$ $6.59\times 10^{-8}$ $6.07\times 10^{-9}$ $39.08\%$ 88 88 $2.22\times$
55 $230$ $8.44\times 10^{-8}$ $7.11\times 10^{-9}$ $29.65\%$ 88 88 $2.85\times$
@@
22 $53$ $6.90\times 10^{-8}$ $6.84\times 10^{-8}$ $2.71\% 44$ 44 $2.18\times$
33 $61$ $1.80\times 10^{-7}$ $1.79\times 10^{-7}$ $21.84\%$ 44 44 $4.14\times$
44 $63$ $1.80\times 10^{-7}$ $1.79\times 10^{-7}$ $12.69\%$ 44 44 $6.87\times$
55 $63$ $1.93\times 10^{-7}$ $1.91\times 10^{-7}$ $8.12\%$ 55 55 $13.59\times$
@@
4.2 Lattice problem
We solve a lattice problem with zero inflow boundary conditions, and its configuration is shown in Fig. 2. The black region corresponds to pure absorption with $(\sigma_{a},\sigma_{s})=(100,0)$, while the white and orange regions correspond to pure scattering with $(\sigma_{a},\sigma_{s})=(0,1)$. A constant source $G(x,y)=1$ is imposed in the orange region. The computational domain $0,5]^{2}$ is discretized using an $80\times 80$ mesh in physical space and the CL(32,16) quadrature in angular space. The scalar fluxes obtained by the low-rank and full-rank SI-DSA are shown on a logarithmic scale in Fig. 2 and match each other well.
This example is used to demonstrate the necessity of using MGS2 with doulbe projection to incrementally update spatial basis and the projected reduced operators. In particular, we compare the performance of our method with direct full QR decomposition and MGS without reprojection. We note that the incremental update of projection operators relies on the hierarchical structure of the spatial basis; therefore, incremental updates of the projected operators cannot be applied when using full QR.
Robustness of the low-rank SI-DSA. As shown in the left plot of Fig. 3, the histories of $\|\bm{\phi}^{(k)}-\bm{\phi}^{(k-1)}\|_{2}$ for the low-rank SI-DSA using MGS2 and the full-rank SI-DSA almost coincide. Similar observations can be made for the convergence histories using MGS and full QR for basis update. As reported in Tab. 2, regardless of using MGS2, MGS or full QR to update the spatial basis, the proposed low-rank SI-DSA always achieves approximately $45.22\%$ compression in DOFs required to store the angular flux, converges in the same number of iterations as the full-rank SI-DSA, and almost the same error with respect to the full-rank reference solution.
Necessity of double projections in MGS and incremental update. Despite the same compression ratio achieved, these three basis update methods differ significantly in their efficiency. As shown in Tab. 2, compared to full-rank SI-DSA, low-rank SI-DSA achieves approximately $1.68\times$ speedup using MGS2, but only $1.15\times$ speedup using MGS, and even becomes slower when using full-QR. This significant performance difference is due to the number of full-QR pass required in the basis update. Using full QR in the basis update of every single inner-loop iteration leads to a total number of $5521$, while MGS reduces this number to $655$ and MGS2 further reduces it to $14$.
Figure 3: Convergence history, effective rank and oversampling ratio during low-rank source iterations for the lattice problem in Sec. 4.2. Left: convergence history. Right: oversampling ratio.
MGS2 $227$ $4.63\times 10^{-7}$ $3.29\times 10^{-8}$ $45.22\%$ $11$ $11$ $1.68\times$
MGS $227$ $4.63\times 10^{-7}$ $3.27\times 10^{-8}$ $45.22\%$ $11$ $11$ $1.15\times$
QR $227$ $4.63\times 10^{-7}$ $3.27\times 10^{-8}$ $45.22\%$ $11$ $11$ $0.65\times$
MGS2 MGS QR
Total number of full QR in the inner-loop basis update $14$ $655$ $5521$
Table 2: Results for the lattice problem in Sec. 4.2. In the basis update step, MGS2, MGS and full QR are applied. MGS2: modified Gram-Schmidt with doulbe projection. MGS: modified Gram-Schimdt. QR: using direct QR decomposition. FR/LR-Iter: number of iterations required for convergence by the full/low-rank SI-DSA.
@@ 4.3 Multiscale variable scattering problem
Robust accuracy and convergence against various $p$ and $q$. Except for $(p,q)=(1,4)$, our low-rank SI-DSA converges in the same number of iterations as the full-rank SI-DSA. When $(p,q)=(1,4)$, it only requires one more iteration. Moreover, the difference between the scalar fluxes produced by the low-rank and full-rank methods, $\|\bm{\phi}_{\textrm{LR}}-\bm{\phi}_{\textrm{FR}}\|_{2}$ is on the order of $10^{-7}$-$10^{-8}$, at least one order of magnitude smaller than the threshold $\epsilon_{\textrm{SI-SA}}=10^{-6}$ used in the stopping criterion of SI-DSA, i.e. $\|\bm{\phi}^{(n)}-\bm{\phi}^{(n-1)}\|_{\infty}\leq\epsilon_{\textrm{SI-SA}}$ .
Robust acceleration and memory saving against $p$ and $q$. We observe that, for various $p$ and $q$, the low-rank SI-DSA consistently reduces the DOFs required to store the angular flux to approximately $38-40\%$ of full rank. Moreover, for $p=1$, slightly more than $2\times$ acceleration is achieved for $q=8,12,16$, and $1.71\times$ for $q=4$. The slight less speedup for $q=4$ is due to the one more outer-loop iteration it requires. For $q=8$, speedup is $2.08\times$ for $p=1$, $1.93\times$ for $p=2$ and $1.43\times$ for $p=3$, consistent with the over-sampling ratio presented in Fig. 5.
Mild space augmentation. In Fig. 5, we present the effective rank and the oversampling ratio during low-rank SI for various $p$ and $q$. With $p=1$, the oversampling ratio remains below $5\%$ for $q=8,12,16$ and below $8\%$ for $q=4$. For fixed $q=8$, increasing the number of sampled angles per iteration leads to a higher oversampling ratio, but it never exceeds $35\%$. These observations verifies the effectiveness of the proposed method in preventing overly aggressive space augmentation during rank adaptation, and implies that a small $p$ value be beneficial to control the oversampling. We also point out that the oversampling ratio in the basis adaptation is $100\%$ for standard augmented BUG [ceruti2022rank] and even higher for standard truncated summation framework. Such oversampling may prevent practical saving in the regime of this example whose effective rank is $38\%$ of the full rank.
@@
44 $292$ $2.95\times 10^{-6}$ $1.35\times 10^{-7}$ $37.64\%$ 77 88 $1.68\times$
88 $296$ $1.53\times 10^{-7}$ $1.34\times 10^{-7}$ $38.16\%$ 77 77 $2.08\times$
$12$ $297$ $1.23\times 10^{-7}$ $4.03\times 10^{-8}$ $38.29\%$ 77 77 $2.04\times$
$16$ $295$ $7.95\times 10^{-8}$ $2.80\times 10^{-8}$ $38.03\%$ 77 77 $1.71\times$
@@
11 $296$ $1.53\times 10^{-7}$ $1.34\times 10^{-7}$ $38.16\%$ 77 77 $2.08\times$
22 $309$ $6.95\times 10^{-7}$ $7.58\times 10^{-8}$ $39.83\%$ 77 77 $1.93\times$
33 $315$ $1.82\times 10^{-6}$ $1.04\times 10^{-7}$ $40.61\%$ 77 77 $1.43\times$
@@ 4.4 Pin-cell problem
Effectiveness and robustness with respect to the randomness in the algorithm. From the accuracy perspective, the difference between the solutions produced by the low-rank SI-DSA and the full-rank SI-DSA is consistently much smaller than $\epsilon_{\textrm{SI-SA}}=10^{-6}$. In all $8$ runs, our low-rank SI-DSA converges with exactly the same number of iterations as its full-rank counterpart.
The solution ranks for all runs of the low-rank SI-DSA are close to $42\%$ of the full rank. Moreover, even in this challenging scenario, thanks to the efficiency of the proposed sampling strategy and the mild space augmentation in the rank-adaptation step, the method achieves at least a $1.88\times$ speedup and an average speedup of $1.96\times$. In the best run, a speedup of approximately $2.01\times$ is obtained.
Figure 7 shows the convergence history and the oversampling ratio during the low-rank SI iterations for the last run. Except for the second-to-last iteration, the change of the scalar flux in the $k$-th iteration, $\|\bm{\phi}^{(k)}-\bm{\phi}^{(k-1)}\|_{2}$, for the low-rank and full-rank SI-DSA almost coincide. The oversampling ratio remains below 8% across all iterations. The observations for the other runs are similar and are therefore omitted.
Finally, we compute the effective rank of the full-rank angular flux by applying truncated SVD to its matrix representation with the same truncation tolerance as in the low-rank method. The effective rank of the full-rank solution is 222, while the ranks of the low-rank solutions range from $209$ to $220$, with an average of 213.50.
@@
Figure 7: Convergence history, effective rank and oversampling ratio during the low-rank SI for the pin-cell problem in Sec. $4.$ Left: convergence history. Right: oversampling ratio.
@@
Solution rank $213.50$ ($\pm 3.34$) $209$ $220$
$\|\psi_{\text{LR}}-\psi_{\text{FR}}\|_{2}$ $1.65(\pm 0.71)\times 10^{-7}$ $7.10\times 10^{-8}$ $3.02\times 10^{-7}$
$\|\phi_{\text{LR}}-\phi_{\text{FR}}\|_{2}$ $8.55(\pm 5.98)\times 10^{-8}$ $1.40\times 10^{-8}$ $1.53\times 10^{-7}$